OpenStudy (anonymous):

given f(x)=1/x and g(x)=x/4, find (f/g)(X)

OpenStudy (anonymous):

Sir, you will devide f(x)/g(x) = (1/x)/(x/4)

OpenStudy (anonymous):

i did, i dont know what to do after that

OpenStudy (kropot72):

\[\frac{1}{x}\div \frac{x}{4}=\frac{1}{x}\times \frac{4}{x}\]

OpenStudy (anonymous):

do i cross multiply?

OpenStudy (kropot72):

Multiply the two numerators to give the numerator of the answer and the two denominators to give the denominator of the answer.